OpenStudy (anonymous):

4√4x^2y * 4√8x^6y^3 4√4x^2y * 4√8x^6y^3 @Mathematics

OpenStudy (anonymous):

Square Root is for which numbers?

OpenStudy (anonymous):

just retype it here using equation editor

OpenStudy (anonymous):

Yes, try using Equation button.

OpenStudy (anonymous):

\[\sqrt[4]{4x^2} * \sqrt[4]{8x^6y^3}\]

OpenStudy (anonymous):

oops in the 4x^2 there's a y after that as well

OpenStudy (anonymous):

ew. to 4th power

OpenStudy (anonymous):

ya it's not fun, my teacher gave this to us for homework but hasn't taught us how to do it :/

OpenStudy (anonymous):

are u sure it isnt cubed root?

OpenStudy (anonymous):

Positive. It's deff 4th root

OpenStudy (anonymous):

\[2x ^{2}\sqrt[4]{2y}\]

OpenStudy (anonymous):

\[8x^{2}y\]

OpenStudy (anonymous):

my answer is wrong. forgot the y

OpenStudy (anonymous):

how do you get your answer? I have a horrible teacher so I'm still a bit lost with the whole root to the power thing :(

OpenStudy (anonymous):

im not sure how abdul got 8... but the answer i got was \[2x^2y \sqrt[4]{2}\]

OpenStudy (anonymous):

How did you get that MJ?

OpenStudy (anonymous):

because they are both to the 4th root, and they are all multiplication you can just multiply whats under the radical and get \[\sqrt[4]{32x^8y^4}\] 32 can be factored into 16*2 16, is 2^4 so you can pull a 2 out and leave a 2 in to get \[2\sqrt[4]{2x^8y^4}\]

OpenStudy (anonymous):

\[\sqrt[4]{x^8}=x^2\] \[\sqrt[4]{y^4}=y\]

OpenStudy (anonymous):

putting it all together gets \[2x^2y \sqrt[4]{2}\]

OpenStudy (anonymous):

Ok, I'm going to trust you MJ lol, I hope you're right :)

OpenStudy (anonymous):

root of 4 can be rewritten as 4^1/2 \[\sqrt[4]{4x^22y*8x^6y^3}\] \[32^{1/4}*x ^{2}*2^{1/4}*y\]

OpenStudy (anonymous):

abdul why is there 32 and a 2?

OpenStudy (anonymous):

4*8 = 32 not 64

OpenStudy (anonymous):

oh wow never mind i cant read

OpenStudy (anonymous):

the real answer is \[4x^2y\] i didnt see that it was 2*y in first problem

OpenStudy (anonymous):

\[64^{1/2}x^2y\]

OpenStudy (anonymous):

wait no i was right the first time ><

OpenStudy (anonymous):

abdul that 2 you were seeing was the x^2 <- that one

OpenStudy (anonymous):

Oh, total mess up, I took that to be 2y. Sorry.. :(

OpenStudy (anonymous):

Oh not it's okay, it's my fault I forgot to put it in the equation :/

OpenStudy (anonymous):

So you guys, if I have \[\sqrt[4]{32x^9}\] am i going to take the 32/ 4?

OpenStudy (anonymous):

What is that power, x^?

OpenStudy (anonymous):

9

OpenStudy (anonymous):

32^1/4*(x^9)^1/4

OpenStudy (anonymous):

you factor 32 in this case 16*2 works the best

OpenStudy (anonymous):

but what do I do with the 9?

OpenStudy (anonymous):

It would become x^9/4

OpenStudy (anonymous):

Ah gottcha!